OpenStudy (anonymous):

Given that p varies directly with q find the constant of variation k if p = -4 and q = 28

OpenStudy (anonymous):

p = k*q -4 = k*28 k = -4/28 k = -1/7

OpenStudy (anonymous):

thankyou very much!